OpenStudy (anonymous):

Simplify the following expression x - 4y + 5z - 3x - y + 2z = ?. 2x - 5y + 7z -2x - 5y - 7z -2x - 5y + 7z -2x -5y + 3z

OpenStudy (cj49):

combine like terms x-3x=-2x -4y-y=-5y 5z+2z=7z -2x-5y+7z

OpenStudy (anonymous):

so is c